OpenStudy (anonymous):

Devon purchased tickets to an air show for 7 adults and 2 children. The total cost was $174 . The cost of a child's ticket was $3 less than the cost of an adult's ticket. Find the price of an adult's ticket and a child's ticket. . adult's ticket: $20; child's ticket: $17 b. adult's ticket: $19; child's ticket: $16 c. adult's ticket: $21; child's ticket: $18 d. adult's ticket: $22; child's ticket: $19

OpenStudy (anonymous):

should have a template for this problem. i have seen it at least 12 times

OpenStudy (wasiqss):

answer is a 20*7+ 2*17=174

OpenStudy (anonymous):

let price of each adult's ticket be x and of each child's ticket be y: from your conditions, 7x+2y=174 -i and y=x-3 -ii now solve eqn i and ii

OpenStudy (anonymous):

substituting eqn ii in eqn i 7x+2(x-3)=174 7x+2x-6=174 9x=180 x=20 from eqn ii y=x-3 y=17 therefore price of adult's ticket is $20 and that of a child is $17